7. Given the initial-value problem y + 3y + 2y = 4x2, y(0) = 3, y0) = 1, a. Find its homogeneous solution using the Consta
Answer #1

7 . using the Given the initial - Value problem y+3y+2y=4m Y(0) = 3, Y(0) = 1 (a) find its homogeneous solution constantAtB n tear therefore the complem mentary Ye (n = qe solution (funetion) -an teek b particular solution of Annihilator methodGeneral solution of the equation is yeu); Ye (M) + Yp (m) Ciec. F +P. I) y you) = (a een tegen + #-64+24²) use the initial co